A 12.0 VV dc battery having no appreciable internal resistance, a 154.0 ΩΩ resistor, an 11.2 mH inductor, and an open switch are all connected in series.

1.After the switch is closed, what is the time constant for this circuit ?

2.After the switch is closed, what is the maximum current that flows through it ?

3.What is the current 73.1 μs after the switch is closed ?

4.After the switch is closed, what is the maximum energy stored in the inductor?
Answer #1

| 0 Tue Constant = % T= 11.2 x 103 154 7.273 x 10 seconds we know (t) = (1-2 ) at t=0 It) = ☆ (1-2) = 0 Hence, current - 0 A